OpenStudy (anonymous):

2^(2x)=8^(x+2) help please

Parth (parthkohli):

2^2x = (2^3)^x+2 2^2x = 2^3x + 6 2x = 3x + 6 -6 = x

OpenStudy (anonymous):

a = 1 b = -2 c = -8 x = - b +/- square root (b^2 - 4ac) / 2a x = 2 +/- square root ( 4 + 32) / 2 = 2 +/- 6 / 2 x1 = -2 x2 = 4

OpenStudy (anonymous):

@ParthKohli is correct.